### v1

# Horizon Entropy Refined: Quantum Contributions and Cosmological Insights

Abstract
We study the effects of quantum fluctuations on the event horizon area and their implications for corrections to the Bekenstein-Hawking entropy. These quantum corrections are incorporated into the framework of large-scale gravitational systems, utilizing the holographic principle to derive modified Friedmann equations. By redefining the Bekenstein-Hawking entropy, our model predicts significant alterations to the Friedmann equations within specific parameter ranges, offering novel perspectives on cosmological scales. Using distance modulus data from the Pantheon supernova sample, we demonstrate the model’s potential to constrain the parameters governing quantum corrections and address unresolved cosmological issues. Crucially, our analysis reveals that quantum fluctuations can increase the area of the event horizon by up to 47%. Beyond this threshold, theoretical predictions encounter substantial challenges when compared with observational data. This approach bridges quantum gravity and observational cosmology, opening new avenues for testing and refining theoretical models.
I Introduction
Exploring the quantum nature of black holes remains a cornerstone in addressing fundamental questions within black hole physics Susskind et al. (2014). While general relativity has traditionally served as the foundation for understanding black holes, quantum effects in the vicinity of the event horizon introduce significant modifications to physical phenomena in these extreme regimes Calmet et al. (2014); Maleki and Maleki (2020).
A particularly striking consequence of the interplay between quantum mechanics and general relativity is the evaporation of black holes through Hawking radiation Hawking (1975). However, these advances have also given rise to profound theoretical challenges, such as the black hole information paradox Bekenstein (1973); Hawking and Ellis (1973); Hooft (1985); Susskind (1997); Callan Jr and Maldacena (1996). Addressing such paradoxes requires a rigorous and careful exploration of quantum phenomena in gravitational settings Susskind et al. (2014); Hawking (2014); Chen et al. (2015); Maleki and Maleki (2022). Moreover, a refined understanding of quantum processes near black holes may pave the way toward a unified framework that seamlessly integrates quantum mechanics with general relativity on cosmological scales.
On the other hand, the discovery of the accelerated expansion of the universe in 1998, driven by Type Ia Supernovae observations Riess et al. (1998), marked a transformative development in cosmology. Subsequent investigations Haridasu et al. (2017) have consistently reaffirmed this acceleration, which, within Einstein’s gravitational framework, is attributed to the cosmological constant $\Lambda$ Zee (2001). The phenomenon driving this accelerated expansion, commonly referred to as Dark Energy (DE), represents one of the most profound mysteries in modern physics. While the cosmological constant provides a straightforward theoretical interpretation of DE, associating it with quantum vacuum energy, the striking discrepancy between theoretical predictions and observed energy scales—known as the cosmological constant problem—remains unresolved Weinberg (2008).
In parallel, the longstanding enigma of Dark Matter (DM) adds another layer of complexity to our understanding of the universe. Initially proposed by Zwicky in 1935 and later substantiated through galactic rotation curves and large-scale structure observations Maleki et al. (2020a); Weinberg (2008), DM plays a pivotal role in the standard cosmological model. The $\Lambda$CDM framework, which combines Cold Dark Matter (CDM) with the cosmological constant, offers a robust explanation for both cosmic acceleration and the gravitational effects of DM. Planck satellite data further elucidate this composition in this model, revealing that the universe consists of approximately 68% dark energy, 28% dark matter, and 4% baryonic matter Aghanim et al. (2018). Despite its successes, the $\Lambda$CDM model encounters notable challenges, as recent observational data reveal inconsistencies that suggest its limitations in fully accounting for certain phenomena. To address these discrepancies, various alternative theories have been proposed, aiming either to refine the $\Lambda$CDM model or to introduce modifications to Einstein’s field equations Weinberg (2008); Maleki et al. (2020b).
A promising avenue for exploring the nature of DE involves the thermodynamic interpretation of gravity, specifically through the holographic principle Susskind (1995); Wang et al. (2017). By incorporating Bekenstein-Hawking entropy into the study of large-scale gravitational systems and utilizing the holographic principle, researchers have successfully derived the Friedmann equations from a thermodynamic perspective Padmanabhan (2014); Wang et al. (2006); Sheykhi (2010, 2018).
This study introduces a modified entropy for Bekenstein-Hawking entropy, derived from quantum fluctuation effects, and examines its implications using cosmological observational data. This modification alters the entropy of the event horizon, resulting in substantial corrections to the derived Friedmann equations. These corrections provide a novel framework for understanding the quantum nature of gravity and its far-reaching impacts on cosmological scales.
II Cosmology of Holographic Entropies
Among various proposals, a significant hypothesis that contributes to understanding the nature of Dark Energy (DE) is based on the holographic principle, known as holographic Dark Energy Susskind (1995); Wang et al. (2017). These models build on the concept that the entropy associated with the boundary is proportional to the area principle, initially proposed by Bekenstein and Hawking in the context of black holes. According to the Bekenstein-Hawking area law, the entropy of a black hole is expressed as Zee (2001)
$S_{\text{BH}}=\frac{A}{4G},$ (1)
where $A$ represents the horizon area, and $G$ denotes the gravitational constant. We also choose units in which Planck’s constant, the speed of light, and the Boltzmann constant are set to unity ($\hslash=c=k_{B}=1$). Thus, through the lens of the gravity-thermodynamics conjecture, we can derive the Friedmann equations from the foundational principles of thermodynamics. It is noteworthy that we introduce the assumption that, once equilibrium is reached, the temperature of the Universe, along with the cosmic horizon, is the same, which is valid in the context of the late-time Universe. Therefore, we can comprehend standard cosmology from an alternative perspective.
However, the crucial issue is related to resolving the crisis faced by the standard model of cosmology. In this direction, akin to modified gravity theories, modifying the entropy associated with the horizon also influences the equations governing the evolution of the universe Nojiri et al. (2021). In fact, there are many proposals for generalized versions of the entropy. Taking into account systems with long-range interactions, such as gravitational systems, which are categorized as non-extensive systems, Tsallis introduced a generalization of standard thermodynamics in 1988 Tsallis (1988). This extension broadens its applicability to non-extensive scenarios while retaining the standard Boltzmann-Gibbs theory as a limiting case Tsallis (1988). It is, thus, noteworthy that, when equipped with a specific entropy expression within any gravity theory, it becomes possible to rephrase the Friedmann equations of the Friedmann-Robertson-Walker (FRW) universe, which are extracted from the first law of thermodynamics on the apparent horizon, and vice versa. Considering this procedure, recent developments have led to significant interest in utilizing non-extensive Tsallis entropy for the entropy of gravitational systems Saridakis et al. (2018); Tavayef et al. (2018); da Silva and Silva (2021). This interest extends to the thermodynamic interpretation of gravitational field equations and the consequential effects of the non-extensive parameter within cosmology. Notably, this can result in variations in the gravitational constant’s strength, which consequently impacts the energy density of the dark components of the universe. This modification can explain the late-time acceleration of the universe without invoking a dark energy component.
Another well-known entropy, which comes from the relativistic generalization of the standard statistical theory, is the so-called Kaniadakis entropy Kaniadakis (2005). This entropy can also be applied to black hole thermodynamics. Using the Kaniadakis entropy, modified Friedmann equations were derived, leading to novel physically significant terms that contribute to an effective dark energy sector of the universe Lymperis et al. (2021). While these theoretical frameworks offer interesting possibilities for exploring the nature of dark energy, benchmarking these models against observational data, such as those from cosmic microwave background radiation or large-scale structure surveys, is essential. The models need to be consistent with the observed behavior of the universe. However, without a reliable theory of quantum gravity, discussing the correction terms to the entropy of a black hole in a precise manner remains elusive.
III Generalized Entropy in Our Model
Here we consider a different approach to suggest an alternative form of entropy, which can be motivated from the quantum gravity perspective.
Figure 1: Surface of a black hole with exaggerated roughness at the Planck scale resolution, illustrating quantum fluctuations and the fuzziness of the event horizon
According to general relativity, the horizon of a black hole is precisely defined by the Schwarzschild radius. On the other hand, from a quantum mechanical perspective, understanding the geometry of spacetime requires measurements using tools such as rods and clocks at different points, each subject to inherent uncertainties. This principle implies that any prospective quantum theory of gravity, built upon the foundations of quantum mechanics, must respect these uncertainty principles.
Therefore, considering the limitations on measurement precision imposed by quantum mechanics, it is reasonable to assume a degree of roughness in the area of a black hole when examining its surface at the Planck scale, as illustrated in Fig. 1. This roughness manifests as a random distribution, potentially rooted in a quantum theory of gravity where uncertainty plays a pivotal role. Evidently, larger surface areas correspond to a greater number of fluctuations, while smaller surfaces exhibit fewer fluctuations.
Assuming that these fluctuations are homogeneously distributed, the number of fluctuations is expected to scale proportionally with the surface area. For instance, doubling the surface area would result in a doubling of the number of fluctuations. Based on this linear relationship, a simple estimate for the corrected area can be expressed as
$\mathcal{A}=A_{0}+\gamma A_{0},$ (2)
where $A_{0}$ represents the area of the black hole corresponding to its Schwarzschild radius, and the second term accounts for the roughness term. Here, $\gamma$ serves as a parameter quantifying the quantum mechanical fuzziness in the black hole’s surface area. To approach the problem from a different perspective that may help enhance our physical intuition, we assume that the black hole surface consists of small pixels of size $l$. We choose this size to be very small so that summing these pixels yields a more accurate representation of the black hole surface. In other words, each pixel contains a few of the smallest possible pixels of size $l_{p}$. Fluctuations in the radius introduce can be of both positive and negative curvature distortions on the surface, each with the size $l_{p}$. Consequently, we can assume that the area of a pixel affected by quantum uncertainty in measurements can be expressed as
$d\mathcal{A}=l^{2}+\eta l_{p}^{2},$ (3)
where $\eta$ is a constant related to the probability distribution of quantum fluctuations. Regardless of whether the surface experiences positive or negative distortion, the area of the selected portion increases by a factor proportional to the Planck area, defined as $A_{p}=l_{p}^{2}$. The larger pixel size is also on the order of a few times the Planck scale (i.e., $l\propto l_{p}$), we can express the correction terms for the black hole’s area as in Eq. 2. It is worth mentioning that, although the volume on average doesn’t change, tiny ripples have the potential to increase the surface by several orders of magnitude. Considering these concepts, we can conceptualize the general form of the black hole entropy as
$S=\frac{A_{0}}{4G}(1+\gamma),$ (4)
In this model, $A_{0}$ once again denotes the horizon area. To test the validity of this model, we explore its cosmological consequences using the thermodynamic interpretation of the gravitational field equations based on the modified entropy. By doing so, we derive the modified Friedmann equations for the universe. This study, in addition to revealing the possible features and constraints of the model, holds the potential to introduce new methods for explaining observational data in cosmology.
IV Deriving the Modified Friedmann Equations
Within the framework of the FRW Universe, the metric’s line element takes the form Zee (2001)
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right],$ (5)
where $a(t)$ is the Universe’s scale factor, $k=0,\pm 1$ denotes the curvature parameter, and $(t,r,\theta,\phi)$ are comoving coordinates. The assumption $a_{0}=a(t=t_{0})=1$ is made, representing the scale factor at the present time.
Taking the apparent horizon as the Universe’s boundary, the associated temperature is determined by Akbar and Cai (2007)
$T_{h}=\frac{-1}{2\pi r_{h}}\left(1-\frac{\dot{r_{h}}}{2Hr_{h}}\right),$ (6)
where $r_{h}={1}/{\sqrt{H^{2}+k/a^{2}}}$ represents the apparent horizon radius. Examining the apparent horizon from a thermodynamic perspective, it emerges as a suitable boundary complying with the first and second laws of thermodynamics. The energy-momentum tensor of the Universe is assumed to take the form $T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $\rho$ denotes the energy density, and $p$ represents the pressure. The conservation law leads to the continuity equation $\dot{\rho}+3H(\rho+p)=0$, where $H={\dot{a}}/{a}$ is the Hubble parameter. The work density associated with the Universe’s expansion is $W={(\rho-p)}/{2}$ Hayward (1998).
To apply the gravity-thermodynamics conjecture, we assert that the first law of thermodynamics on the apparent horizon holds
$dE=T_{h}dS+WdV.$ (7)
Denoting the total energy of the universe as $E=\rho V$, where $V=\frac{4}{3}\pi r_{h}^{3}$, differentiating this expression and substituting $\dot{\rho}$ from the continuity equation, we obtain
$dE=4\pi r_{h}^{2}\rho dr_{h}-4\pi Hr_{h}^{3}(\rho+p)dt.$ (8)
From the entropy relation in Eq.4, we find $dS=2\pi(1+\gamma)r_{h}dr_{h}/G$. Considering the horizon temperature and applying the first law of thermodynamics, we get
$\frac{1+\gamma}{4\pi Gr_{h}^{3}}dr_{h}=H(\rho+p)dt.$ (9)
Using the continuity equation and integrating both sides, we derive
$\rho=\frac{3(1+\gamma)}{8\pi Gr_{h}^{2}}.$ (10)
Considering the relation between $r_{h}$ and the cosmological parameters, we can rewrite Eq.10 as
$\left(H^{2}+\frac{k}{a^{2}}\right)=\frac{8\pi G}{3(1+\gamma)}\rho.$ (11)
This equation provides a generalized form of the first Friedmann equation, where the modified factor $(1+\gamma)$ comes from the entropy correction.
The second Friedmann equation can be obtained by differentiating Eq.11 and combining it with the continuity equation. After some algebra, we arrive at
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3(1+\gamma)}(\rho+3p).$ (12)
It is worth mentioning that although the root of the correction comes from the quantum mechanical fuzziness in the area of the black hole, it is possible to consider these effects as changing the gravitational constant, such that
$G_{\text{eff}}=\frac{G}{1+\gamma}.$ (13)
Figure 2: Evolution of the scale factor for various parameter choices of generalized entropy in a flat, matter-dominated universe. The solid line represents the scenario without corrections to the Bekenstein-Hawking entropy. The figure illustrates the effects of parameter on evolution of the universe.
To analyze the model, we first derive the relevant cosmological parameters. For a flat, matter-dominated universe, using Eq.11, the scale factor is expressed as
$a(t)=\left(\frac{6\pi G}{1+\gamma}\rho_{0}\right)^{1/3}t^{2/3},$ (14)
The scale factor is modified by a factor of ${1}/{(1+\gamma)^{1/3}}$. In Fig. 2, we illustrate the evolution of the scale factor for a flat, matter-dominated universe under five distinct parameter choices. The solid line represents the scenario without any corrections to the entropy, corresponding to the Bekenstein-Hawking entropy. In this case, the universe rapidly enters a decelerating phase. To address this challenge, the introduction of a dark energy component becomes necessary.
Intriguingly, altering the parameters of the generalized entropy modifies this behavior, producing results that align more closer with observational data in some range of parameters. The age of the universe can be computed using Eq.14, which gives
$t_{\text{age}}=\frac{2}{3}\sqrt{1+\gamma}\,t_{H_{0}},$ (15)
Here, $t_{H_{0}}=1/H_{0}$ denotes the Hubble time. Deviations from the Bekenstein-Hawking entropy are accounted for through the correction parameter $\gamma$ in Eq.4. As illustrated in Fig.2, these adjustments have a substantial impact on the evolution of the scale factor and the universe’s expansion dynamics.
From Fig.2, it can be observed that for $1.20<\gamma<$ 1.2, the model effectively modifies standard cosmological predictions. However, for $\gamma$ values exceeding this range, estimating the age of the universe becomes problematic. This analysis imposes a constraint on the permissible range of the correction term $\gamma$, ensuring consistency with observational data.
To deepen our understanding of entropy in this context, we analyze the luminosity distance and its dependence on redshift, both of which are pivotal for studying cosmic dynamics. The luminosity distance, which quantifies how an object’s apparent brightness relates to its intrinsic luminosity, plays a central role in probing the universe’s expansion. It is a function of redshift ($z$), which measures the stretching of light caused by cosmic expansion. The relationship is expressed as:
$d_{L}(z)=(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})},$ (16)
This expression encapsulates how the changing expansion rate impacts the observed brightness of distant objects, revealing key insights into the universe’s acceleration and the influence of dark energy.
In our case, we can use the modified form of the first Friedmann equation (Eq.11) for a flat universe and apply the scale factor-redshift relation. Substituting the Hubble parameter $H(z)=H_{0}(1+z)^{3/2}/(\sqrt{1+\gamma})$, we can insert this into Eq.16. After performing the integration, the result is
$d_{L}(z)=\frac{2\sqrt{1+\gamma}}{H_{0}}\left(1+z-\sqrt{1+z}\right)$ (17)
which due to the entropy modification, the supernova appears in larger distances by a factor of $\sqrt{1+\gamma}$.
In cosmology, the distance modulus ($\mu$) is a widely used parameter, defined as $\mu\equiv m-M$, where $m$ represents the apparent brightness and $M$ is the absolute brightness of an object. Its relationship with the luminosity distance is expressed as
$\mu=5\log\left(\frac{d_{L}}{\text{Mpc}}\right)+25,$ (18)
where $d_{L}$ is the luminosity distance in Megaparsecs (Mpc).
In Fig. 3, we present the distance modulus as a function of redshift for 1,048 supernovae from the Pantheon sample Scolnic et al. (2018) and the predictions of our model in a flat, matter-only universe. The dot-dashed line represents the matter-only universe without considering the modified entropy. The supernovae in the dataset appear fainter (or more distant) than predicted by the matter-only model. The solid line represents the best fit to the data using the modified entropy model with $\gamma=0.21$, while the dashed line corresponds to the case of $\gamma=1.0$, which diverges much from the observational data.
Figure 3: Distance modulus ($\mu$) versus redshift of 1048 supernovae from the Pantheon sample Scolnic et al. (2018). The solid line represents the best-fit with $\gamma=0.21$ for the modified entropy, while the dashed line corresponds to a flat matter-only universe. The supernovae clearly appear fainter (or more distant) than predicted in a matter-only universe.
To evaluate the deviation of the model from observational data and determine the best fit, we employed the chi-squared ($\chi^{2}$) parameter, defined as
$\chi^{2}=\sum_{i=1}^{N}\frac{(O_{i}-M_{i})^{2}}{\sigma_{i}^{2}},$ (19)
Where $O_{i}$ is the $i$-th observed value, $M_{i}$ is the $i$-th model prediction, $\sigma_{i}$ is the uncertainty in the $i$-th observation and $N$ is the total number of data points. The $\chi^{2}$ statistic provides a measure of how well the model fits the data. A smaller value of $\chi^{2}$ indicates a better fit to the observations. In Fig. 4, we plot the normalized $\chi_{n}^{2}$, which is the $\chi^{2}$ value divided by the degrees of freedom, defined as the total number of data points minus the number of free parameters. In our case, this value is $1047$.
Figure 4: Normalized $\chi^{2}$ ($\chi_{n}^{2}$) as a function of the roughness parameter $\gamma$. The plot highlights the improvement in the fit at $\gamma=0.21$. The shaded gray region indicates the range where the model operates without significant observational inconsistencies.
We observe that for $\gamma=0.21$, the model fits the data significantly better and can effectively account for the role of dark energy to a good extent. However, for more precise results, contributions from dark energy-like models, in addition to matter, remain necessary. For instance, if we compute the deceleration factor, it becomes $0.5$, which indicates that the matter-only universe in this model decelerates as expected. However, due to the weakness of gravity in this model, the universe expands more, showing potential to reduce the gap between a matter-only universe and observational reality. Furthermore, from the Fig. 4, we deduce that for $\gamma>0.47$, this model becomes problematic in cosmology. Interestingly, the negative effects of quantum fluctuations in the area of the event horizon also present challenges, supporting the intuitive picture discussed in Sec. III. This analysis consequently imposes constraints on $\gamma$, the quantum mechanical correction parameter. The shaded gray region represents the parameter range where the model performs reliably, avoiding major observational discrepancies. As a result, the area of the event horizon can increase by at most 47% due to quantum fluctuations. Beyond this threshold, new challenges emerge when attempting to reconcile theoretical predictions with observational data.
V Summary and Conclusion
We have investigated the impact of quantum corrections to the event horizon entropy on cosmological dynamics, providing new insights into the interplay between quantum gravity and large-scale structure. By incorporating quantum fluctuations into the black hole surface area, the classical Bekenstein-Hawking entropy is generalized with a correction term proportional to a roughness parameter, $\gamma$. According to the thermodynamics-gravity conjecture, any correction to the entropy leads to modifications of the Friedmann equations, effectively introducing a $\gamma$-dependent factor that alters the evolution of the scale factor, $a(t)$, and the universe’s expansion history.
Our analysis reveals that small corrections produce negligible deviations from standard cosmology, while moderate values yield significant changes in cosmological behavior. Larger corrections ($\gamma>0.47$) result in inconsistencies with observational constraints, including the distance modulus versus redshift data from the supernovae data set. This analysis of compatibility with empirical data demonstrates that quantum fluctuations can lead to a maximum $47\%$ increase in event horizon area without conflicting with observations. These findings highlight the sensitivity of the universe’s evolution to quantum-gravitational effects and offer an alternative framework to address cosmological problems. The study provides a quantitative basis for further exploration of quantum gravity influence on cosmology, with observational constraints on $\gamma$ serving as a critical bridge between theory and experiment. Future work could refine these constraints with different observational datasets and extend the framework to include additional quantum-gravitational phenomena. This investigation underlines the potential of quantum entropy corrections to address unresolved questions in fundamental physics and cosmology.

### v2

@@
 Notably, this can result in variations in the gravitational constant’s strength, which consequently impacts the energy density of the dark components of the universe.
-This modification can explain the late-time acceleration of the universe without invoking a dark energy component.
+This modification can explain the late-time acceleration of the universe without invoking a dark energy component Tavayef et al.
+(2018); Zamora and Tsallis (2022).
 Another well-known entropy, which comes from the relativistic generalization of the standard statistical theory, is the so-called Kaniadakis entropy Kaniadakis (2005).
@@
 III Generalized Entropy in Our Model
-Here we consider a different approach to suggest an alternative form of entropy, which can be motivated from the quantum gravity perspective.
-Figure 1: Surface of a black hole with exaggerated roughness at the Planck scale resolution, illustrating quantum fluctuations and the fuzziness of the event horizon
+In this section, we propose an alternative form of entropy inspired by considerations from quantum gravity.
 According to general relativity, the horizon of a black hole is precisely defined by the Schwarzschild radius.
-On the other hand, from a quantum mechanical perspective, understanding the geometry of spacetime requires measurements using tools such as rods and clocks at different points, each subject to inherent uncertainties.
-This principle implies that any prospective quantum theory of gravity, built upon the foundations of quantum mechanics, must respect these uncertainty principles.
-Therefore, considering the limitations on measurement precision imposed by quantum mechanics, it is reasonable to assume a degree of roughness in the area of a black hole when examining its surface at the Planck scale, as illustrated in Fig.
+However, from a quantum mechanical perspective, understanding the geometry of spacetime requires measurements involving tools such as rods and clocks.
+These measurements are subject to inherent uncertainties Garay (1995); Kempf et al.
+(1995).
+For instance, attempting to measure a distance with high precision, especially at very short scales, encounters limitations imposed by the Heisenberg uncertainty principle.
+Precisely determining the position of the measuring rod introduces significant uncertainty in its momentum, which in turn results in uncertainty in the rod’s energy.
+Since energy contributes to spacetime curvature through general relativity, this energy uncertainty induces distortions in spacetime, leading to a fluctuating geometry.
+Thus, measurements at extremely small scales inherently face precision limits due to the interplay of quantum mechanics and general relativity.
+This phenomenon persists even in the weak-energy regime of quantum gravity.
+Any consistent theory that fully reconciles gravity with quantum mechanics must accommodate this fundamental argument.
+The interplay of quantum uncertainty, energy fluctuations, and spacetime curvature highlights a fundamental limitation in measurement precision, underscoring the need for a deeper understanding of quantum spacetime.
+Figure 1: An exaggerated representation of a black hole’s surface at Planck-scale resolution, illustrating quantum fluctuations and the fuzziness of the event horizon.
+Given the limitations on measurement precision imposed by quantum mechanics, it is reasonable to assume a degree of roughness in the surface area of a black hole when examined at the small scales, as shown in Fig.
 1.
-This roughness manifests as a random distribution, potentially rooted in a quantum theory of gravity where uncertainty plays a pivotal role.
-Evidently, larger surface areas correspond to a greater number of fluctuations, while smaller surfaces exhibit fewer fluctuations.
-Assuming that these fluctuations are homogeneously distributed, the number of fluctuations is expected to scale proportionally with the surface area.
-For instance, doubling the surface area would result in a doubling of the number of fluctuations.
-Based on this linear relationship, a simple estimate for the corrected area can be expressed as
+This roughness arises as a random distribution, potentially rooted in a quantum theory of gravity where uncertainty plays a pivotal role.
+Larger surface areas correspond to more fluctuations, while smaller surfaces exhibit fewer fluctuations.
+Assuming that these fluctuations are homogeneously distributed, the number of fluctuations scales proportionally with the surface area.
+For example, doubling the surface area results in a doubling of the number of fluctuations.
+Based on this proportionality, the corrected surface area can be expressed as:
 $\mathcal{A}=A_{0}+\gamma A_{0},$ (2)
-where $A_{0}$ represents the area of the black hole corresponding to its Schwarzschild radius, and the second term accounts for the roughness term.
-Here, $\gamma$ serves as a parameter quantifying the quantum mechanical fuzziness in the black hole’s surface area.
-To approach the problem from a different perspective that may help enhance our physical intuition, we assume that the black hole surface consists of small pixels of size $l$.
-We choose this size to be very small so that summing these pixels yields a more accurate representation of the black hole surface.
-In other words, each pixel contains a few of the smallest possible pixels of size $l_{p}$.
-Fluctuations in the radius introduce can be of both positive and negative curvature distortions on the surface, each with the size $l_{p}$.
-Consequently, we can assume that the area of a pixel affected by quantum uncertainty in measurements can be expressed as
-$d\mathcal{A}=l^{2}+\eta l_{p}^{2},$ (3)
-where $\eta$ is a constant related to the probability distribution of quantum fluctuations.
-Regardless of whether the surface experiences positive or negative distortion, the area of the selected portion increases by a factor proportional to the Planck area, defined as $A_{p}=l_{p}^{2}$.
-The larger pixel size is also on the order of a few times the Planck scale (i.e., $l\propto l_{p}$), we can express the correction terms for the black hole’s area as in Eq.
-2.
-It is worth mentioning that, although the volume on average doesn’t change, tiny ripples have the potential to increase the surface by several orders of magnitude.
-Considering these concepts, we can conceptualize the general form of the black hole entropy as
+where $A_{0}$ represents the area of the black hole as defined by its Schwarzschild radius, and the second term accounts for the surface roughness.
+Here, $\gamma$ is a parameter quantifying the quantum mechanical fuzziness in the black hole’s surface area.
+While the average volume remains unchanged, tiny ripples can significantly increase the surface area and intuitively we expect $\gamma$ is always positive.
+Building on this concept, the generalized form of black hole entropy can be written as:
 $S=\frac{A_{0}}{4G}(1+\gamma),$ (3)
-In this model, $A_{0}$ once again denotes the horizon area.
-To test the validity of this model, we explore its cosmological consequences using the thermodynamic interpretation of the gravitational field equations based on the modified entropy.
-By doing so, we derive the modified Friedmann equations for the universe.
-This study, in addition to revealing the possible features and constraints of the model, holds the potential to introduce new methods for explaining observational data in cosmology.
+where $A_{0}$ again denotes the horizon area.
+To test the validity of this model, we examine its cosmological implications by employing the thermodynamic interpretation of the gravitational field equations, modified by the generalized entropy.
+This leads to the derivation of modified Friedmann equations for the universe.
+By exploring these modifications, we can uncover new features and constraints of the model and potentially introduce innovative methods for explaining observational data in cosmology.
 IV Deriving the Modified Friedmann Equations
@@
 It is a function of redshift ($z$), which measures the stretching of light caused by cosmic expansion.
-The relationship is expressed as:
+The relationship is expressed as Weinberg (2008)
 $d_{L}(z)=(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})},$ (15)
@@
 In cosmology, the distance modulus ($\mu$) is a widely used parameter, defined as $\mu\equiv m-M$, where $m$ represents the apparent brightness and $M$ is the absolute brightness of an object.
-Its relationship with the luminosity distance is expressed as
+Its relationship with the luminosity distance is expressed as Weinberg (2008)
 $\mu=5\log\left(\frac{d_{L}}{\text{Mpc}}\right)+25,$ (17)
@@
 The supernovae clearly appear fainter (or more distant) than predicted in a matter-only universe.
-To evaluate the deviation of the model from observational data and determine the best fit, we employed the chi-squared ($\chi^{2}$) parameter, defined as
+To evaluate the deviation of the model from observational data and determine the best fit, we employed the chi-squared ($\chi^{2}$) parameter, defined as Witkov and Zengel (2019)
 $\chi^{2}=\sum_{i=1}^{N}\frac{(O_{i}-M_{i})^{2}}{\sigma_{i}^{2}},$ (18)
@@
 The study provides a quantitative basis for further exploration of quantum gravity influence on cosmology, with observational constraints on $\gamma$ serving as a critical bridge between theory and experiment.
+Moreover, this model logically follows from the effects of quantum mechanics’ random distributions on spacetime geometry in the low-energy limit.
+Consequently, any quantum gravity theory consistent with quantum mechanical foundations is expected to exhibit similar properties.
 Future work could refine these constraints with different observational datasets and extend the framework to include additional quantum-gravitational phenomena.